OpenStudy (kathy_morales13):

can someone please help me? A cheetah can travel up to 29 meters per second. what is the cheetahs speed,to the nearest tenth,in miles per hour ?

OpenStudy (wholock):

Okay, so if it can travel 29 meters per second, how many meters can it travel in a minute? 29 * 60 = ?

OpenStudy (kathy_morales13):

1,740

OpenStudy (wholock):

Alright, sorry had a bit of a brain fail there, to make it easier, there are 3600 seconds in a hour, so 29 * 3600 = please forgive me

OpenStudy (wholock):

Now, 1 mile = 1609 meters. so it is a proportion problem.

OpenStudy (wholock):

29 * 3600 = 104, 400 Now to convert meters to miles, divide the meters by 1609 104,400/1609 = 64.9 miles

OpenStudy (wholock):

To round to the nearest tenth, simply round to 65. The cheetah can run 65 mph

OpenStudy (kathy_morales13):

got it thank you :)

OpenStudy (wholock):

Your welcome!